1/12 + 2 /3 + 4/9=
Answer by stanbon(75887)

1/12 + 2 /3 + 4/9=
====
lcd = 36
=====
Rewrite each fraction with the lcd as its denominator:
3/lcd + (2*12)/lcd + (4*4)/lcd
-----
Combine over the lcd:
[3+24+16}/lcd
-----
= 44/(36
----
= 11/9
---
= 1 2/9
==========================
Cheers,
Stan H.